OpenStudy (anonymous):

Solve -20 = 5/6x X= I think that X is -24 @alexandervonHumboldt

OpenStudy (studybuddy8):

Yep

OpenStudy (anonymous):

x does = -24

OpenStudy (anonymous):

help with a few more please

OpenStudy (anonymous):

Evaluate x - 4 / 2 when x = -1/6. -1.6 - 4 / 2 = -1.6 - 2 Answer =

OpenStudy (anonymous):

i think that Answer is -3.6

OpenStudy (studybuddy8):

yes! your on a role :)

OpenStudy (anonymous):

yes -20=5/6x 6/5* -20=6/5* 5/6x -120/5=x -24=x